THE MIDPOINT OF A LINE SEGMENT AB IS THE POINT (-1, 2).THE X COORDINATE OF A IS 5 AND THE Y-COORDINATE OF B IS -9,FIND THE POINTS OF A+B
Answer by Menjax(62)

First off please don't use caps to gain attention it's unnecessary...
using formulas
Mx=(x1+x2)/2
My=(y1+y2)/2
-1=(5+x2)/2
2=(y1-9)/2
Point A is at (5,-5)
Point B is at (-7,-9)